You are required to standardize a dilute solution of sulphuric acid, and undertake the following steps: (i) 4.805 g of solid sodium tetraborate (Na2B4O7) are added to a 250 mL volumetric flask, and the solution made up to the mark with distilled water. (ii) 25.00 mL aliquots of the stock solution from (i) are transferred to conical flasks, for titration against the sulphuric acid solution in the burette. (iii) You perform titrations using two different indicators: bromocresol green, and phenolphthalein. With bromocresol green, the average titre volume is 22.50 cm3 , while with phenolphthalein, the average volume is 9.50 cm3 . (a) Write the balanced equation for the reaction between Na2B4O7 and sulphuric acid


1
Expert's answer
2021-07-19T04:04:57-0400

 On adding a calculated quantity of concentrated sulphuric acid to a hot concentrated solution of borax, boric acid is produced.


Borax ( Na2B4O7 ) + H2SO4 = Boric Acid (H3BO3 ) + Na2SO4 .


Molar mass of Na2B4O7 = 201 .22 g / mol .

Given mass of Na2B4O7 = 4.805 g .

moles of Na2B4O7 = 4.805 / 201.22 = 0.024 mol .

Volume = 250 ml .= 0.25 litre .


Concentration ( molarity in mol per litre ) = 0.024 / 0.25 = 0.096 M .


Milli Moles of Na2B4O7 = 25 * 0.096 = 2.4 milli moles .